OpenStudy (anonymous):

Identify the vertex for the graph of y = 3x^2 + 12x + 5. (2, 41) (2, 35) (-2, -31) (-2, -7)

OpenStudy (anonymous):

use -b/2a to find x then put the value of x in equation and find y

OpenStudy (anonymous):

b=12 a=3

OpenStudy (anonymous):

so its -12/2(3) =-12/6 =-2

OpenStudy (anonymous):

x=-2 now y will be 3(-2)^2+12(-2)+5

OpenStudy (anonymous):

which is 12-24+5 =-7

OpenStudy (anonymous):

so answer is (-2,-7) :)